a 10.0 liter container holds one of the noble gases, but which one is unknown. The mass of the gas is 25.25 g, while the pressure is 3.05 atm and the temperature is 24 C. Calculate the molar mass and identify the gas


1
Expert's answer
2021-04-27T07:48:45-0400

According to the ideal gas law:

pV = nRT = mRT/Mr

where p - pressure, V - volume, n - number of moles, R - gas constant (R = 0.082 L⋅atm⋅K−1⋅mol−1), T - temperature (T = 24 °C = 297.15 K), m - mass, Mr - molar mass.

From here:

Mr = mRT / pV = 25.25 g × 0.082 L⋅atm⋅K−1⋅mol−1 × 297.15 K / (3.05 atm × 10.0 L) = 20.2 g/mol

As a result, the unknown gas is neon.


Answer: 20.2 g/mol, neon